OpenStudy (anonymous):

A radial saw has a blade with a 7-inch radius. Suppose that the blade spins at 1300 rpm. (radians per minute). a.) find the angular speed of the blade in rad/min. b.) find the linear speed of the saw teeth in ft/s (feet per second). USE (3.1416 for pi)

OpenStudy (anonymous):

ok no, it did not specify that rpm meant radians per second

OpenStudy (anonymous):

in part "a" it says radians per minute so I didn't know if that corresponded to the question

ganeshie8 (ganeshie8):

I'll assume 1300 rpm = 1300 revolutions per minute,

ganeshie8 (ganeshie8):

then, for part a) : angular speed = \(2 \pi f = 2\pi*1300 = 2600 \pi \) rad/min

ganeshie8 (ganeshie8):

for pat b) : linear speed = \(r \omega \) \(r\) = radius \(\omega\) = angular speed

ganeshie8 (ganeshie8):

you will need to convert the units .. .

OpenStudy (anonymous):

i have never seen these formulas before lol

ganeshie8 (ganeshie8):

u will see it more... if you're doing uniform-circular motion problems.... :)

OpenStudy (anonymous):

okay is that all i will need to do? ima keep this question open in case something comes up

OpenStudy (anonymous):

what does the long f mean, after 2 pi?

ganeshie8 (ganeshie8):

f = frequency = 1300 revolutions per minute

OpenStudy (anonymous):

okay lol thanks

ganeshie8 (ganeshie8):

linear speed = \(\large r \omega = 7 \times 2600 \pi \) inches/ minute

ganeshie8 (ganeshie8):

convert it to feet/second

OpenStudy (anonymous):

so 7/12 for inches to feet and...8168(2600pi) would be 490080 in seconds?

ganeshie8 (ganeshie8):

1/12 for inches 1/60 for seconds

ganeshie8 (ganeshie8):

careful, minute is in

ganeshie8 (ganeshie8):

careful, minute is in denomunator

ganeshie8 (ganeshie8):

\(\large \mathbb{7\times 2600 \pi \frac{inch}{min} }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{7\times 2600 \pi \frac{12inch}{12min} }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{7\times 2600 \pi \frac{feet}{12min} }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{7\times 2600 \pi \frac{feet}{12\times 60 seconds} }\)

ganeshie8 (ganeshie8):

\(\large \mathbb{\frac{7\times 2600 \pi}{12\times 60} \frac{feet}{ seconds} }\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

okay thank you so much!!!

ganeshie8 (ganeshie8):

you should get 79.4 feet/second

ganeshie8 (ganeshie8):

u wlc :) btw, u may use wolfram for calculator : http://www.wolframalpha.com/input/?i=%287*+2600+%5Cpi%29%2F%2812%5Ctimes+60%29

OpenStudy (anonymous):

wow that is an awesome website thanks for all the help!

ganeshie8 (ganeshie8):

yes it is.... np =)